OpenStudy (anonymous):

the velocity of a particle moving along he x-axis for time t(greater than or equal to)0 is given by v(t)=sin t m/sec. .......

OpenStudy (anonymous):

what is the particles displacement from t=0 to t=3Pi/2 seconds?

OpenStudy (anonymous):

it is the integral of sint from t=0 to t=3pi/2

OpenStudy (anonymous):

let s be the displacement, then \[s = \int\limits_{0}^{3\pi/2}\sin t dt= - \cos t _{0}^{3\pi/2}=-(0-1)= 1 m\]

OpenStudy (anonymous):

oh wow! Thank you so much! math is just not my thing even though we went over these since the beginnign of time!

OpenStudy (anonymous):

you're welocme :) is this a course in calculus?

OpenStudy (anonymous):

yes. it's ap calculus at my high school. but honestly my teacher teaches it way too fast because he is obsessed with us passing the ap exam. i think he would be better off teaching at a college level.

OpenStudy (anonymous):

what would the distance the particle traveled during the time interval be?

OpenStudy (anonymous):

do you first know the difference between the displacement and distance?!

OpenStudy (anonymous):

yes displacement is like the amount you have taken overall. say your were at 9 than you took 3 steps left your displacement would be 12, right?

OpenStudy (anonymous):

let's assume you were is position (in the number line) 2 and then you move two steps to the right and then 1 step to the left..so you're now in position 3. okay?

OpenStudy (anonymous):

your displacement in this case is l3-2l=1

OpenStudy (anonymous):

so it is just where you are. not how many steps you have taken...

OpenStudy (anonymous):

but the total distance you traveled is 3, which is 2 right steps and 1 left step

OpenStudy (anonymous):

right ok. got it. thank you. are you like a math major or professor or just a math super genious?

OpenStudy (anonymous):

lol none of them :P

OpenStudy (anonymous):

than what are you?

OpenStudy (anonymous):

if you notice the displacement can be calculated by taking the what the distance you walk to the right minus the distance you walk to the left..( which is 2 -1 =1) while the total distance is just the sum of them (2+1=3)

OpenStudy (anonymous):

a human being I guess

OpenStudy (anonymous):

oh ok... you cant be in high school...to know all that..

OpenStudy (anonymous):

I am a university student.

OpenStudy (anonymous):

hmmm. well. i can not be positive but i'm coming to the conlusion that you are smart. at least when it comes to math. :)

OpenStudy (anonymous):

haha that could be right

OpenStudy (anonymous):

lol. well even though i am so not done with my math work and need tons more help, i must go to my next class because the bell rang... can not wait until this summer to start college so that i can pick my own times to go to school! lol. Thanks for all your help. it wont be long before i am back on though :)

OpenStudy (anonymous):

Have fun :) I miss high school by the way!!